### v2

# Higher Gaussian maps on the hyperelliptic locus and second fundamental form

Abstract.
In this paper we study higher even Gaussian maps of the canonical bundle on hyperelliptic curves and we determine their rank, giving explicit descriptions of their kernels. Then we use this descriptions to investigate the hyperelliptic Torelli map $j_{h}$ and its second fundamental form. We study isotropic subspaces of the tangent space $T_{{\mathcal{H}}_{g},[C]}$ to the moduli space ${\mathcal{H}}_{g}$ of hyperelliptic curves of genus $g$ at a point $[C]$, with respect to the second fundamental form $\rho_{HE}$ of $j_{h}$. In particular, for any Weierstrass point $p\in C$, we construct a subspace $V_{p}$ of dimension $\lfloor\frac{g}{2}\rfloor$ of $T_{{\mathcal{H}}_{g},[C]}$ generated by higher Schiffer variations at $p$, such that the only isotropic tangent direction $\zeta\in V_{p}$ for the image of $\rho_{HE}$ is the standard Schiffer variation $\xi_{p}$ at the Weierstrass point $p\in C$.
D. Faro, P. Frediani and A. Lacopo are members of GNSAGA (INdAM) and are partially supported by PRIN project Moduli spaces and special varieties (2022).
1. Introduction
The purpose of this paper is twofold. On the one hand we study higher even Gaussian maps of the canonical bundle on hyperelliptic curves and we determine their rank, giving an explicit description of their kernels. On the other hand, we use this description to investigate the local geometry of the image via the Torelli map of the moduli space ${\mathcal{H}}_{g}$ of hyperelliptic curves of genus $g$ in the moduli space ${\mathcal{A}}_{g}$ of principally polarised abelian varieties of dimension $g$.
Let $C$ be a hyperelliptic curve of genus $g\geq 3$. Denote by $I_{2}$ the kernel of the multiplication map
$\mu_{0}:S^{2}H^{0}(C,K_{C})\rightarrow H^{0}(C,K_{C}^{\otimes 2})^{+},$
where $H^{0}(C,K_{C}^{\otimes 2})^{+}$ denotes the invariant subspace of $H^{0}(C,K_{C}^{\otimes 2})$ under the action of the hyperelliptic involution. The canonical map is the composition of the $2:1$ map $\pi:C\rightarrow{\mathbb{P}}^{1}$ with the $(g-1)-{th}$ Veronese embedding $\nu_{g-1}:{\mathbb{P}}^{1}\rightarrow{\mathbb{P}}^{g-1}$, hence the space $I_{2}$ is identified with the vector space of quadrics containing the rational normal curve $\nu_{g-1}({\mathbb{P}}^{1})$ in ${\mathbb{P}}^{g-1}$. The second Gaussian map $\mu_{2}$ is a linear map
$\mu_{2}:I_{2}=Ker(\mu_{0})\rightarrow H^{0}(C,K_{C}^{\otimes 4}),$
and more generally, for any $k\geq 1$, the $2k$-th Gaussian map is a linear map
$\mu_{2k}:Ker(\mu_{2k-2})\rightarrow H^{0}(C,K_{C}^{\otimes 2k+2}).$
These maps were introduced by Wahl in [18]. One can also define odd Gaussian maps, each one being defined on the kernel of the previous one. The most important and studied one is the first Wahl map:
$\mu_{1}:\wedge^{2}H^{0}(C,K_{C})\rightarrow H^{0}(C,K_{C}^{\otimes 3}).$
It was proven by Wahl that if $C$ sits on a $K3$ surface, then the first Wahl map $\mu_{1}$ is not surjective (see [18], see also [1]). On the other hand, Ciliberto, Harris and Miranda proved that for the general curve of genus $g\geq 10$, $g\neq 11$, $\mu_{1}$ is surjective (see [2], [16] for another proof).
If $C$ is a hyperelliptic curve, it was proven in [17], [3] that the rank of $\mu_{1}$ is $32g-$ 3, while in [4] it was shown that the rank of $\mu_{2}$ is $52g-$ 5.
Our first main result is the following
Theorem 1.1.
(See Theorem 3.1) Let $C$ be a hyperelliptic curve of genus $g\geq 3$ Then for every $0\leq k\leq\frac{g-1}{2}$
(1) $Rank(\mu_{2k})=2g-(4k+1),$
(2) $dim(Ker(\mu_{2k}))=\frac{(g-1)(g-2)}{2}-k(2g-2k-3).$
Then, $Rank(\mu_{2k})=0$ for every $k>\lfloor\frac{g-1}{2}\rfloor$.
From this we immediately see (Remark 3.2) that if $g$ is odd we have the following chain of inclusions:
$0=Ker(\mu_{g-1})\subsetneq Ker(\mu_{g-3})\subsetneq...\subsetneq Ker(\mu_{2})\subsetneq I_{2};$
while if $g$ is even we have
$0=Ker(\mu_{g-2})\subsetneq Ker(\mu_{g-4})\subsetneq...\subsetneq Ker(\mu_{2})\subsetneq I_{2}.$
Denote by ${\mathcal{H}}_{g}$ the hyperelliptic locus in ${\mathcal{M}}_{g}$ and by
$j_{h}:{\mathcal{H}}_{g}\rightarrow{\mathcal{A}}_{g}$
the restriction of the Torelli map to ${\mathcal{H}}_{g}$.
The map $j_{h}$ is an orbifold immersion (see [15]). The main tool for studying the local geometry of the hyperelliptic Torelli locus $j_{h}({\mathcal{H}}_{g})\subset{\mathcal{A}}_{g}$ is the second fundamental form of the map $j_{h}$. Recall that the moduli space ${\mathcal{A}}_{g}$ is the quotient of the Siegel space $Sp(2g,{\mathbb{R}})/U(g)$ under the action of $Sp(2g,{\mathbb{Z}})$. We endow ${\mathcal{A}}_{g}$ with the Siegel metric, that is the orbifold metric induced by the symmetric metric on the Siegel space. The second fundamental form of $j_{h}$ is a map
$\rho_{HE}:N^{*}_{\mathcal{H}_{g}|\mathcal{A}_{g}}\rightarrow Sym^{2}\Omega^{1}_{\mathcal{H}_{g}},$
where $N^{*}_{\mathcal{H}_{g}|\mathcal{A}_{g}}$ denotes the conormal bundle of $j_{h}(\mathcal{H}_{g})$ in $\mathcal{A}_{g}$. At a point $[C]\in{\mathcal{H}}_{g}$, $\rho_{HE}$ is therefore a linear map
$\rho_{HE}:I_{2}\rightarrow Sym^{2}(H^{0}(C,K_{C}^{\otimes 2})^{+}).$
One of the main problems in the study of the local geometry of the image of the hyperelliptic locus in ${\mathcal{A}}_{g}$ under the Torelli map is to investigate the existence of totally geodesic subvarieties of ${\mathcal{A}}_{g}$ contained in the hyperelliptic locus.
This is related with the hyperelliptic Coleman-Oort conjecture, that says that for $g\geq 8$ there do not exist positive dimensional special subvarieties of ${\mathcal{A}_{g}}$ generically contained in the hyperelliptic Torelli locus, namely contained in the closure of $j_{h}({\mathcal{H}}_{g})$ and intersecting $j_{h}({\mathcal{H}}_{g})$. This conjecture was recently proven by Moonen [14], generalising a previous result of Lu and Zuo [13].
In fact, special subvarieties are totally geodesic, hence the study of the second fundamental form is related with the above problem.
Here we take a different viewpoint, namely, we are interested in the behaviour of the second fundamental form of the map $j_{h}$ in relation with higher even Gaussian maps.
In [5] it was proven that the map $\rho_{HE}$ can be expressed in terms of the Hodge Gaussian map $\rho$ introduced in [8]. An important property proven in [8] is a formula computing, for any $Q\in I_{2}$, the Hodge Gaussian map $\rho(Q)$ on Schiffer variations at points $p$ on any curve $C$ in terms of the second Gaussian map $\mu_{2}(Q)$.
This property is one main tool used in [11, Theorem 6.2] to show that the maximal dimension of a germ of a totally geodesic subvariety of ${\mathcal{A}}_{g}$ contained in the hyperelliptic locus is $g+1$.
In the case of hyperelliptic curves, the result of [8] gives the following expression for $\rho_{HE}(Q)$ on the Schiffer variation $\xi_{p}\in H^{1}(C,T_{C})^{+}$ at a Weierstrass point $p\in C$, in terms of the second Gaussian map :
$\rho_{HE}(Q)(\xi_{p}\odot\xi_{p})=-2\pi i\mu_{2}(Q)(p).$
Then we have $\rho_{HE}(Q)(\xi_{p}\odot\xi_{p})=0$, since $p$ is a Weierstrass point and one can easily see that for all the quadrics $Q\in I_{2}$, we have $\mu_{2}(Q)(p)=0$ (see Remark 6.13).
In [9], a relation between higher even Gaussian maps $\mu_{2k}$ of the canonical bundle on a smooth projective curve of genus $g\geq 4$ and the Hodge Gaussian map $\rho$ is given, generalising the above result.
Here we exploit similar techniques to compute $\rho_{HE}(Q)$ for quadrics $Q$ contained in the kernel of higher Gaussian maps on odd higher Schiffer variations $\xi_{p}^{2k+1}\in H^{1}(C,T_{C})^{+}$ at a Weierstrass point $p\in C$ (see Section 4 for the definition of higher Schiffer variations). For any $k\leq 2g-3$, consider the $(k+1)$-dimensional subspace
$V_{k}:=\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{2k+1}\rangle\subset H^{1}(T_{C})^{+}.$
We prove the following
Theorem 1.2.
(See Theorems 6.6, 6.7).
  1. (1)
    Let $0\leq k\leq\lfloor\frac{g-3}{2}\rfloor$ and let $Q\in Ker(\mu_{2k})$. Then we have
    $\rho_{HE}(Q)(\xi_{p}^{l}\odot\xi_{p}^{m})=0\quad\forall l+m\leq 4k+3,\ l,m\ odd.$
    This implies that the subspace $V_{k}=\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{2k+1}\rangle\subset H^{1}(T_{C})^{+}$ is isotropic for $\rho_{HE}(Q)$, $\forall Q\in Ker(\mu_{2k}).$
  2. (2)
    There exists $Q\in Ker(\mu_{2k})$ such that $\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+3})\neq 0$. Hence the subspace $V_{k+1}=\langle\xi_{p}^{1},...,\xi_{p}^{2k+1},\xi_{p}^{2k+3}\rangle\subset H^{1}(T_{C})^{+}$ is not isotropic for $\rho_{HE}(Ker(\mu_{2k}))$.
Following [7] we call a nonzero direction $\zeta\in H^{1}(T_{C})^{+}$ asymptotic if
$\rho_{HE}(Q)(\zeta\odot\zeta)=0,\ \forall Q\in I_{2}.$
Clearly tangent directions to totally geodesic subvarieties are asymptotic directions. Notice that saying that a nonzero element $\zeta\in H^{1}(T_{C})^{+}$ is an asymptotic direction means that the point $[\zeta]\in{\mathbb{P}}H^{1}(T_{C})^{+}$ is in the base locus of the linear space of quadrics $\rho_{HE}(I_{2})\subset Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$. In [5] it is shown that $\rho_{HE}$ is injective, hence the image of $\rho_{HE}$ in $Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$ has dimension $\frac{(g-1)(g-2)}{2}$. So $\rho_{HE}(I_{2})$ is a linear system of quadrics in ${\mathbb{P}}(H^{1}(C,T_{C})^{+})\cong{\mathbb{P}}^{2g-2}$ of dimension $\frac{(g-1)(g-2)}{2}$.
Thus if $g\leq 5$, there are asymptotic directions, in fact there are also examples of special subvarieties of ${\mathcal{A}}_{g}$ of positive dimension generically contained in the hyperelliptic locus (see e.g. [10]).
On the other hand, for $g$ sufficiently high one would expect that the intersection of a space of quadrics of dimension $\frac{(g-1)(g-2)}{2}$ in ${\mathbb{P}}^{2g-2}$ would be empty, hence that asymptotic directions would not exist. This is not true because Schiffer variations at Weierstrass points are asymptotic directions (see Remark 6.13). On the other hand we prove the following
Theorem 1.3.
(See Theorem 6.14).
Let $C$ be a hyperelliptic curve of genus $g\geq 3$ and $p\in C$ be a Weierstrass point. Then the asymptotic directions in the space $V_{\lfloor\frac{g-2}{2}\rfloor}$ are exactly the Schiffer variations $\xi_{p}=\xi_{p}^{1}$.
As a consequence we have the following
Remark 1.4.
(see Remark 6.15).
In [5, Corollary 3.4] it is shown that $\rho_{HE}$ is injective, so
$\dim\rho_{HE}(Ker\mu_{2k})=\dim Ker\mu_{2k}=\frac{(g-1)(g-2)}{2}-k(2g-2k-3),$
for all $k$. For any hyperelliptic curve $C$ of genus $g\geq 3$ and for any Weiestrass point $p\in C$, we have proven that we have the following chain of subspaces of quadrics in ${\mathbb{P}}H^{1}(T_{C})^{+}$ and of corresponding maximal isotropic subspaces in $V_{\frac{g-3}{2}}$ if $g$ is odd and in $V_{\frac{g-2}{2}}$ if $g$ is even.
If $g$ is odd:
$\rho_{HE}(Ker(\mu_{g-3}))\subsetneq\rho_{HE}(Ker(\mu_{g-5}))...\subsetneq\rho_{HE}(Ker(\mu_{2}))\subsetneq\rho_{HE}(I_{2});$
$V_{\frac{g-3}{2}}\supsetneq V_{\frac{g-5}{2}}\supsetneq...\supsetneq V_{1}\supsetneq V_{0}=\langle\xi_{p}\rangle,$
while if $g$ is even we have
$\rho_{HE}(Ker(\mu_{g-4}))\subsetneq\rho_{HE}(Ker(\mu_{g-6}))...\subsetneq\rho_{HE}(Ker(\mu_{2}))\subsetneq\rho_{HE}(I_{2}),$
$V_{\frac{g-4}{2}}\supsetneq V_{\frac{g-6}{2}}\supsetneq...\supsetneq V_{1}\supsetneq V_{0}=\langle\xi_{p}\rangle.$
Finally, we show that Theorem 6.14 also allows us to deduce Corollary 6.16, which gives a bound for the dimension of a germ of a totally geodesic submanifold of ${\mathcal{A}}_{g}$, generically contained in the Hyperelliptic Torelli locus. Nevertheless this is weaker than the one already proven in [11, Theorem 6.2].
The structure of the paper is as follows.
In Section 2 we recall the definition and basic properties of Gaussian maps. In Section 3 we determine the kernel of all even higher Gaussian maps on the hyperelliptic locus and we compute their rank, proving Theorem 1.1. In Section 4 we define and give the basic properties of higher Schiffer variations. In Section 5 we recall the definition and some results on the second fundamental form of the hyperelliptic Torelli map $j_{h}$. Finally in Section 6 we prove Theorems 1.2, 1.3 and their consequences (Remark 6.15 and Corollary 6.16).
2. Gaussian maps
Let $C$ be a smooth curve of genus $g$, and let $L$ and $M$ be two line bundles on $C$. In this section we recall the definition of the (higher) Gaussian maps on $C$ associated with the line bundles $L$ and $M$.
Set $S:=C\times C$ and let $\Delta\subset S$ be the diagonal. Take a non-negative integer $k$, then the $k$-th Gaussian (or Wahl map) associated with $L$ and $M$ is the map given by restriction to the diagonal
$H^{0}(S,L\boxtimes M(-k\Delta))\stackrel{{\scriptstyle\Phi_{k,L,M}}}{{\longrightarrow}}H^{0}(S,L\boxtimes M(-k\Delta)_{|{\Delta}})\cong H^{0}(C,L\otimes M\otimes K_{C}^{\otimes k}).$
We will only consider the case $L=M$, and we set $\Phi_{k,L}:=\Phi_{k,L,L}$. Since we have the identification $H^{0}(S,L\boxtimes M)\cong H^{0}(C,L)\otimes H^{0}(C,M)$, the map $\Phi_{0,L}$ is given by the multiplication map of global sections
$H^{0}(C,L)\otimes H^{0}(C,L)\rightarrow H^{0}(C,L^{\otimes 2}),$
which vanishes identically on $\wedge^{2}H^{0}(C,L)$. Then we have
$\ker\Phi_{0,L}=H^{0}(S,L\boxtimes L(-\Delta))\cong\wedge^{2}H^{0}(C,L)\oplus I_{2}(L),$
where $I_{2}(L)$ is the kernel of the multiplication map $S^{2}H^{0}(C,L)\rightarrow H^{0}(C,L^{\otimes 2})$. Since $\Phi_{1,L}$ vanishes on symmetric tensors, one writes
$\mu_{1,L}:={\Phi_{1,L}}_{|\wedge^{2}H^{0}(L)}:\wedge^{2}H^{0}(L)\rightarrow H^{0}(K_{C}\otimes L^{\otimes 2}).$
Take a local frame $l$ for $L$, local coordinate $z$ and two sections $s_{1},s_{2}\in H^{0}(C,L)$ with $s_{i}=f_{i}(z)l$. Then we have
(3) $\mu_{1,L}(s_{1}\wedge s_{2})=(f^{\prime}_{1}f_{2}-f^{\prime}_{2}f_{1})dz\otimes l^{2}.$
Consequently, the zero divisor of $\mu_{1,L}(s_{1}\wedge s_{2})$ is $2B+R$, where $B$ is the base locus of the pencil $\langle s_{1},s_{2}\rangle$ and $R$ is the ramification divisor of the associated morphism.
Again $H^{0}(S,L\boxtimes L(-2\Delta))$ decomposes as the sum of $I_{2}(L)$ and the kernel of $\mu_{1,L}$. Since $\Phi_{2,L}$ vanishes on skew-symmetric tensors, we write
$\mu_{2,L}:={\Phi_{2,L}}_{|I_{2}(L)}:I_{2}(L)\rightarrow H^{0}(C,L^{\otimes 2}\otimes K_{C}^{\otimes 2}).$
We will denote by $\mu_{2}$ the second gaussian map of the canonical line bundle $K_{C}$ on $C$:
$\mu_{2}:=\mu_{2,K_{C}}:I_{2}(K_{C})\rightarrow H^{0}(K_{C}^{\otimes 4}).$
If $\alpha=\sum l_{i}\otimes m_{i}\in I_{2}(L)$ with $l_{i}=f_{i}(z)l$ and $m_{i}=g_{i}(z)l$, using that $\mu_{0,L}(\alpha)=0$, and $\Phi_{1,L}(\alpha)=0$, one finds the identities
(4) $\sum f_{i}g_{i}\equiv 0,\ \ \sum f_{i}^{\prime}g_{i}-f_{i}g_{i}^{\prime}\equiv 0.$
Taking derivatives of the first equation in (4) and using the second one we obtain the identities:
(5) $\sum f_{i}^{\prime}g_{i}\equiv\sum f_{i}g_{i}^{\prime}\equiv 0,$
Derivating the equations in (5) we have:
(6) $\sum f_{i}^{\prime\prime}g_{i}=-\sum f_{i}^{\prime}g_{i}^{\prime}=\sum f_{i}g_{i}^{\prime\prime}.$
The local expression of $\mu_{2,L}$ is
(7) $\mu_{2,L}(\alpha)=\sum f_{i}^{\prime\prime}g_{i}\ dz^{\otimes 2}\otimes l^{\otimes 2}=\sum f_{i}g_{i}^{\prime\prime}\ dz^{\otimes 2}\otimes l^{\otimes 2}=-\sum f_{i}^{\prime}g_{i}^{\prime}dz^{\otimes 2}\otimes l^{\otimes 2}.$
More generally the local expression for $\Phi_{2k,L}$ when $k\geq 1$ is given by
(8) $\Phi_{2k,L}(\alpha)={(-1)^{k}}\sum f_{i}^{(k)}g_{i}^{(k)}\ dz^{\otimes 2k}\otimes l^{\otimes 2},$
and since it vanishes on antisymmetric tensors one defines
(9) $\mu_{2k,L}:=\Phi_{2k,L_{|_{Ker(\mu_{2k-2,L})}}}.$
Analogously the local expression for $\Phi_{2k+1,L}$ when $k\geq 0$ is given by
(10) $\Phi_{2k+1,L}(\alpha)=\sum(f_{i}^{(k+1)}g_{i}^{(k)}-f_{i}^{(k)}g_{i}^{(k+1)})\ dz^{\otimes 2k+1}\otimes l^{\otimes 2},$
and since it vanishes on symmetric tensors one defines for every $k\geq 2$
(11) $\mu_{2k+1,L}:=\Phi_{2k+1,L_{|_{Ker(\mu_{2k-1,L})}}}.$
Moreover since $\alpha=\sum l_{i}\otimes m_{i}\in Ker(\mu_{2k-2,L})$ if and only if
(12) $\mu_{0,L}(\alpha)=\Phi_{0,L}(\alpha)=...=\mu_{2k-2,L}(\alpha)=\Phi_{2k-1,L}(\alpha)=0,$
and these conditions give identities
(13) $\sum f_{i}^{(h)}g_{i}^{(r)}\equiv 0,\forall h,r\ \text{such that}\ h+r\leq 2k-1,$
the local expression for $\mu_{2k,L}$ is equivalent to
(14) $\mu_{2k,L}(\alpha)=(-1)^{m}\sum f_{i}^{(2k-m)}g_{i}^{(m)}dz^{\otimes 2k}\otimes l^{\otimes 2},$
for every $m=0,...,2k$. Analogously $\alpha=\sum l_{i}\otimes m_{i}\in Ker(\mu_{2k-1,L})$ if and only if
(15) $\Phi_{0,L}(\alpha)=\mu_{1,L}(\alpha)=...=\mu_{2k-1,L}(\alpha)=\Phi_{2k,L}(\alpha)=0,$
and these conditions give the identities
(16) $\sum f_{i}^{(h)}g_{i}^{(r)}\equiv 0,\forall h,r\ \text{such that}\ h+r\leq 2k,$
and from this, it follows that we can equivalently express (10) as
(17) $\mu_{2k+1,L}(\alpha)=(-1)^{m}\sum(f_{i}^{(2k+1-m)}g_{i}^{(m)}-f_{i}^{(m)}g_{i}^{(2k+1-m)})dz^{\otimes 2k+1}\otimes l^{\otimes 2}$
for every $m=0,...,k$.
Remark 2.1.
Observe that for every $k\geq 1$
(18) $Ker(\Phi_{2k,L})=Ker(\mu_{2k,L})\oplus Ker(\mu_{2k-1,L}),$
and from the definition of $\mu_{k,L}$ we have the inclusions:
(19) ... $\subset Ker(\mu_{2k,L})\subset Ker(\mu_{2(k-1),L})\subset...\subset Ker(\mu_{2,L})\subset I_{2}=Ker(\mu_{0,L});$
(20) ... $\subset Ker(\mu_{2k+1,L})\subset Ker(\mu_{2k-1,L})\subset...\subset Ker(\mu_{1,L}).$
3. Rank of higher order Gaussian maps on hyperelliptic curves
Let $C$ be a hyperelliptic curve of genus $g\geq 3$, $L$ be the line bundle giving the $g^{1}_{2}$. Set $M=K_{C}\otimes L^{\vee}$ and $\pi:C\rightarrow{\mathbb{P}}^{1}$ the map induced by $|L|$. Denote by $\nu_{n}:{\mathbb{P}}^{1}\hookrightarrow{\mathbb{P}}^{n}$ the $n^{th}$ Veronese embedding. Recall that the canonical map is given by the composition $\nu_{g-1}\circ\pi$, so
$K_{C}\cong\pi^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(g-1))\cong L^{\otimes(g-1)}$
and $M\cong L^{\otimes(g-2)}$. Then $H^{0}(C,M)\cong H^{0}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}(g-2))$ has dimension $g-1$.
Denote by $\Phi_{k}$, $\mu_{k}$, the Gaussian maps $\Phi_{k,K_{C}}$ and $\mu_{k,K_{C}}$. The main result of this section is the computation of the rank of higher Gaussian maps for any hyperelliptic curve of any genus. More precisely we are going to prove the following:
Theorem 3.1.
Let $C$ be a hyperelliptic curve of genus $g\geq 3$ Then for every $0\leq k\leq\frac{g-1}{2}$
(21) $Rank(\mu_{2k})=2g-(4k+1),$
(22) $dim(Ker(\mu_{2k}))=\frac{(g-1)(g-2)}{2}-k(2g-2k-3).$
Then, $Rank(\mu_{2k})=0$ for every $k>\lfloor\frac{g-1}{2}\rfloor$.
Notice that the result was already known for $\mu_{2}$ (see [4]).
Remark 3.2.
By Theorem 3.1, it is immediate to see that if $g$ is odd we have the following chain of inclusions:
$0=Ker(\mu_{g-1})\subsetneq Ker(\mu_{g-3})\subsetneq...\subsetneq Ker(\mu_{2})\subsetneq I_{2}(K_{C});$
while if $g$ is even we have
$0=Ker(\mu_{g-2})\subsetneq Ker(\mu_{g-4})\subsetneq...\subsetneq Ker(\mu_{2})\subsetneq I_{2}(K_{C}).$
Notice that if $g$ is odd, $\dim(Ker(\mu_{g-3}))=1$, and if $g$ is even $\dim(Ker(\mu_{g-4}))=3$.
Let us first recall the following well known result (see for example [4]).
Lemma 3.3.
Let $C$ be a hyperelliptic curve of genus $g\geq 3$. Let $|L|$ be the $g^{1}_{2}$. Set $M=K_{C}\otimes L^{\vee}$, let $\omega_{1},...,\omega_{g-1}$ be a basis for $H^{0}(M)$ and let $\langle s,t\rangle$ be a basis for $H^{0}(L)$. Then the map defined by
[…]
$\omega_{i}\wedge\omega_{j}$ $\rightarrow Q_{ij}:=s\omega_{i}\odot t\omega_{j}-s\omega_{j}\odot t\omega_{i},$
is an isomorphism. In particular, observe that $\{Q_{ij}\}$ gives a basis for $I_{2}$.
In [4, Lemma 4.1] it is shown the following
Lemma 3.4.
For any $Q_{ij}$ as in the previous lemma:
$\mu_{2}(Q_{ij})=\mu_{1,L}(s\wedge t)\mu_{1,M}(\psi^{-1}(Q_{ij})),$
where with the expression
(23) $\mu_{1,L}(s\wedge t)\mu_{1,M}(\psi^{-1}(Q_{ij}))$
we mean the image of $\mu_{1,L}(s\wedge t)\otimes\mu_{1,M}(\psi^{-1}(Q_{ij}))$ under the multiplication map
(24) $H^{0}(K_{C}\otimes L^{\otimes 2})\otimes H^{0}(K_{C}\otimes M^{\otimes 2})\rightarrow H^{0}(K_{C}^{\otimes 4}).$
Observe that from Lemma 3.4 it follows that $Rank(\mu_{2})=Rank(\mu_{1,M})$ since $\mu_{1,L}(s\wedge t)$ is a nonzero section in $H^{0}(K_{C}\otimes L^{\otimes 2})$. Indeed the zero locus of $\mu_{1,L}(s\wedge t)$ is given by the base locus of $|L|$ together with the ramification divisor of the induced morphism (recall (3)).
We start generalizing Lemma 3.4 to any higher-order Gaussian maps. We use the same notations as in Lemma 3.3
Lemma 3.5.
Let $k\geq 0$ be an integer and let
$Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}Q_{ij}\in Ker(\mu_{2k}).$
Then
  • (i)
    for any $k\geq 1$
    $\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j})\in Ker(\mu_{2k-1,M}),$
  • (ii)
    for any $k\geq 0$
    $\mu_{2k+2}(Q)=(k+1)\mu_{1,L}(s\wedge t)\mu_{2k+1,M}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j}))\in H^{0}(K_{C}^{\otimes{2k+4}}),$
Proof.
Let us proceed by induction. When $k=0$ $(ii)$ is Lemma 3.4, and when $k=1$ $(i)$ it is an immediate consequence of the hypothesis $\mu_{2}(Q)=0$ together with Lemma 3.4.
Now take $n\geq 2$ and suppose that $(ii)$ holds for every $10\leq k<n-$ 1 and $(i)$ holds for every $1\leq k<n$. We are going to prove that $(ii)$ holds for $k=n-1$, which automatically implies that $(i)$ holds for $k=n$. Set $k=n-1$ and suppose that
$Q=\sum\limits_{1\leq 1<j\leq g-1}a_{ij}Q_{ij}\in Ker(\mu_{2k}).$
On some open sets write
$\omega_{i}=f_{i}dz^{\prime},\ t=fT\ \text{and}\ s=gT,$
where $f_{i}$ is a holomorphic function, $dz^{\prime}$ is a local generator on $M=K_{C}\otimes L^{\vee}$ and $T$ is a local generator on $L$. From the definition of $Q_{ij}$ in Lemma 3.3, it follows that we can write locally $Q$ as
(25) $\sum\limits_{1\leq 1<j\leq g-1}a_{ij}Q_{ij}=\sum a_{ij}(ff_{i}dz\odot gf_{j}dz-ff_{j}dz\odot gf_{i}dz),$
where we can take $z$ to be a local coordinate on $C$. Then, by definition
$\mu_{2k+2}(\sum\limits_{1\leq 1<j\leq g-1}a_{ij}Q_{ij})$
(26) $=\sum\limits_{1\leq 1<j\leq g-1}a_{ij}((ff_{i})^{(k+1)}(gf_{j})^{(k+1)}-(ff_{j})^{(k+1)}(gf_{i})^{k+1})dz^{\otimes(2k+4)}.$
Observe that the latter expression can be written as
(27) $=\left[\sum\limits_{1\leq i<j\leq g-1}a_{ij}\sum\limits_{h=0}^{k+1}\binom{k+1}{h}f^{(k+1-h)}f_{i}^{(h)}\sum\limits_{l=0}^{k+1}\binom{k+1}{l}g^{(k+1-l)}f_{j}^{(l)}\right]dz^{\otimes(2k+4)}$
$-\left[\sum\limits_{1\leq i<j\leq g-1}a_{ij}\sum\limits_{r=0}^{k+1}\binom{k+1}{h}f^{(k+1-h)}f_{j}^{(h)}\sum\limits_{e=0}^{k+1}\binom{k+1}{l}g^{(k+1-l)}f_{i}^{(l)}\right]dz^{\otimes(2k+4)}$
(28) $=\left[\sum\limits_{h,l=0}^{k+1}\binom{k+1}{h}\binom{k+1}{l}f^{(k+1-h)}g^{(k+1-l)}\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(h)}f_{j}^{(l)}-f_{j}^{(h)}f_{i}^{(l)})\right]dz^{\otimes(2k+4)}$
By hypothesis $Q\in Ker(\mu_{2k})$ and hence from the inductive hypothesis
$\mu_{2k-1,M}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j}))=0.$
Then using (16) and (3) it follows that for every $h+l\leq 2k$
$\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(h)}f_{j}^{(l)}-f_{j}^{(h)}f_{i}^{(l)})\equiv 0.$
Then (28) becomes
$\left[\binom{k+1}{k}fg^{(1)}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(k+1)}f_{j}^{(k)}-f_{j}^{(k+1)}f_{i}^{(k)})\right]dz^{\otimes(2k+4)}+$
$\left[\binom{k+1}{k}f^{(1)}g(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(k)}f_{j}^{(k+1)}-f_{j}^{(k)}f_{i}^{(k+1)})\right]dz^{\otimes(2k+4)}.$
This is just
$\left[\binom{k+1}{k}(fg^{(1)}-f^{(1)}g)(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(k+1)}f_{j}^{(k)}-f_{j}^{(k+1)}f_{i}^{(k)})\right]dz^{\otimes(2k+4)},$
which is equal to
$(k+1)\mu_{1,L}(t\wedge s)\mu_{2k+1,M}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j})),$
and so $(ii)$ holds. ∎
In the following lemma, we are going to describe the equations of the loci
$\subset Ker(\mu_{2k})\subset Ker(\mu_{2(k-1)})\subset...\subset Ker(\mu_{2})\subset I_{2}=Ker(\mu_{0}).$
Consider a hyperelliptic curve given by the affine equation:
(29) $y^{2}=\prod_{i=1}^{2g+2}(x-t_{i}).$
For simplicity from now on we will assume $t_{1}=0$, so $|M|=|K_{C}(-2p)|$, where $p$ is the Weierstrass point $(0,0)$ in affine coordinates. Consider the basis
(30) $\{\alpha_{0}=\frac{dx}{y},\alpha_{1}=\frac{xdx}{y},...,\alpha_{g-1}=\frac{x^{g-1}dx}{y}\}$
of $H^{0}(K_{C})$. Then a basis of $H^{0}(M)$ is given by
(31) $\{\alpha_{1}=\frac{xdx}{y},...,\alpha_{g-1}=\frac{x^{g-1}dx}{y}\}.$
The equations of $Ker(\mu_{2})$ in $I_{2}$ in terms of this basis have already been described in the proof of [4, Proposition 4.2]. Indeed we have the following
Lemma 3.6.
Let
$Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}Q_{ij}\in I_{2},$
$Q_{ij}$ as in 3.3, with respect to the basis of $H^{0}(M)$ given by $\alpha_{i}=x^{i}\frac{dx}{y}$, for $i=1,...,g-1$. Then $Q\in Ker(\mu_{2})$ if and only if for all $33\leq l\leq 2g-$ 3
(32) $\sum\limits_{\begin{subarray}{c}1\leq i<j\\ i+j=l\end{subarray}}a_{ij}(j-i)=0.$
Now we generalize the approach in [4]. The strategy is to use Lemma 3.5 together with the explicit description of the basis $\alpha_{0},...,\alpha_{g-1}$ of the $H^{0}(C,K_{C})$ for a hyperelliptic curve $C$. We have the following result.
Lemma 3.7.
Let
$Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}Q_{ij}\in I_{2},$
with $Q_{ij}=s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i}$ as in 3.3, for $i=1,...,g-1.$ For any $k\geq 2$, $Q\in Ker(\mu_{2k})$ if and only if
  • $\forall\ 3\leq l\leq 2g-3$,
    $\sum\limits_{\begin{subarray}{c}1\leq i<j\\ i+j=l\end{subarray}}a_{ij}(j-i)=0,$
  • $\forall\ 2\leq m\leq k$, $\forall\ 2m-1\leq l\leq 2g-3$,
    (33) $\sum\limits_{\begin{subarray}{c}1\leq i<j\\ i+j=l\\ i\geq m-1,\\ j\geq m-1\end{subarray}}a_{ij}(j-i)ij(i-1)(j-1)...(i-(m-2))(j-(m-2))=0.$
Proof.
We will argue locally. Take the basis of $H^{0}(C,K_{C})$ given in (30) and the one for $H^{0}(C,M)$ given in (31). We will proceed by induction on $k$. The thesis holds for $k=1$ by Lemma 3.6. Now we suppose that the thesis holds for $k-1$, $k\geq 2$ and take $Q\in Ker(\mu_{2k-2})$. Since $Q$ also belongs to the previous kernels, proving the lemma is equivalent to prove that $Q\in Ker(\mu_{2k})$ if and only if
(34) $\sum\limits_{\begin{subarray}{c}1\leq i<j\\ i+j=l\\ i\geq k-1,\\ j\geq k-1\end{subarray}}a_{ij}(j-i)ij(i-1)(j-1)...(i-(k-2))(j-(k-2))=0,$
for all $32k-1\leq l\leq 2g-$ 3. Using lemma 3.5 we have that $\mu_{2k}(Q)=0$ if and only if $\mu_{2k-1,L}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j}))=0$, that is if and only if
(35) $\sum\limits_{1\leq i<j\leq g-1}a_{ij}((x^{i})^{(k-1)}(x^{j})^{(k)}-(x^{i})^{(k)}(x^{j})^{(k-1)})\equiv 0,$
where we are taking $x$ as local coordinate. Observe that (35) is equal to
(36) $\sum\limits_{\begin{subarray}{c}1\leq i<j\leq g-1\\ i\geq k-1\\ j\geq k\end{subarray}}a_{ij}i...(i-(k-2))x^{(i-(k-1))}j...(j-(k-1))x^{j-k}$
(37) $-\sum\limits_{\begin{subarray}{c}1\leq i<j\leq g-1\\ i\geq k\\ j\geq k-1\end{subarray}}a_{ij}i...(i-(k-1))x^{(i-k)}j...(j-(k-2))x^{j-(k-1)},$
which can be written after some simple algebraic manipulations as
(38) $\sum\limits_{\begin{subarray}{c}1\leq i<j\leq g-1\\ i\geq k-1\\ j\geq k-1\end{subarray}}x^{(i+j-(2k-1))}a_{ij}i...(i-(k-2))j...(j-(k-2))(j-i).$
So we have shown that for $Q\in Ker\mu_{2k-2}$, $\mu_{2k}(Q)=0$ if and only if for every $32k-1\leq l\leq 2g-$ 3
$\sum\limits_{\begin{subarray}{c}1\leq i<j\leq g-1\\ i+j=l\\ i\geq k-1\\ j\geq k-1\end{subarray}}a_{ij}(j-i)i...(i-(k-2))j...(j-(k-2))=0.$
This concludes the proof. ∎
Now we come to the proof Theorem 3.1. The strategy of the proof is clear: we have to determine the number of linearly independent equations in (33).
Proof of Theorem 3.1.
We will start introducing some notations. For every $i,j$ such that $11\leq i<j\leq g-$ 1, for every $l=3,...,2g-3$, let $(w_{l})^{2}$ the vector whose coordinates are
$(w_{l})^{2}_{ij}=\begin{cases}(j-i)\ \ \text{if}\ i+j=l\\ 0\ \ \text{if}\ i+j\neq l\end{cases}$
which are the coefficients of the $l$-th equation of the set of equations given in Lemma 3.7, describing $Ker(\mu_{2})$, ordered by increasing values of $i$. For every $2\leq r\leq k$ and $l=2r-1,...,2g-3$, let $w_{l}^{(2r)}$ be the vector whose coordinates are
(39) $(w_{l})^{(2r)}_{ij}=\begin{cases}(j-i)ij(i-1)(j-1)...(i-(r-2))(j-(r-2))\ \ \text{if}\ i+j=l,i,j\geq r-1\\ 0\ \ \text{if}\ i+j\neq l,\ \text{or}\ 1\leq i\leq r-2.\end{cases}$
which are the coefficients of the $l$-th equation of the set of equations which describe $Ker(\mu_{2r})$ inside $(Ker(\mu_{2(r-1)}))$, ordered by increasing values of $i$.
For any $33\leq l\leq 2g-$ 3 set
(40) $n_{l}=\#\{(i,j):1\leq i<j\leq g-1,\ i+j=l\}.$
and let $B^{\prime}_{k,l}$ be the $k\times n_{l}$ matrix whose $r$th row, $1\leq r\leq k$, are the $n_{l}$ coordinates of $w_{l}^{(2r)}$ corresponding to the indexes $(i,j)$ such that $i+j=l$.
Observe that $B^{\prime}_{k,l}$ is just the matrix with rows $w_{l}^{(2)},...,w_{l}^{(2k)}$ where we have removed the entries corresponding to $i+j\neq l$. Notice that these entries are all 00 for all the vectors $w_{l}^{(2r)}$.
Define $c_{k,l}=min\{n_{l},k\}$ and let $B_{k,l}$ be the minor of $B^{\prime}_{k,l}$ of order $c_{k,l}$ given by the first $c_{k,l}$ rows and $c_{k,l}$ columns. We are going to prove that it is not zero for all $33\leq l\leq 2g-$ 3 by induction on $k$.
Consider first the case $k=2$. Observe that if $l=3,4,2g-4,2g-3$, then $n_{l}=1$. In this case $B_{2,l}=j-i>0$ where $(i,j)$ is the only pair such that $i+j=l$. If $55\leq l\leq 2g-$ 5 then $n_{l}\geq 2$ and $c_{2,l}=min(2,n_{l})=2$ and we have
$B_{2,l}=\begin{pmatrix}j-i&j-i-2\\ ij(j-i)&(j-i-2)(i+1)(j-1)\end{pmatrix}$
where $i$ is the minimum such that $i+j=l$. Observe that
$det(B_{2,l})=(j-i)(j-i-2)det\begin{pmatrix}1&1\\ ij&(i+1)(j-1)\end{pmatrix}$
Set
$A_{2,l}:=\begin{pmatrix}1&1\\ ij&(i+1)(j-1)\end{pmatrix}$
Subtracting the first column of $A_{2,l}$ from the second we obtain the matrix
$\begin{pmatrix}1&0\\ ij&j-i-1\end{pmatrix}$
whose determinant is $j-i-1>0$. Now assume that for every $2\leq k_{0}\leq k$ and for every $33\leq l\leq 2g-$ 3, $B_{k,l}$ is not zero. We want to prove that the same holds for $k+1$. If $n_{l}\leq k$, then $n_{l}<k+1$, $c_{k+1,l}=min\{n_{l},k\}=n_{l}$ and $B_{k+1,l}=B_{k,l}$. In this case, we have nothing to prove. Assume then $n_{l}>k$. Then $c_{k+1,l}=min\{n_{l},k+1\}=k+1$. Then $B_{k+1,l}$ is the following $(k+1)\times(k+1)$ matrix
$\begin{pmatrix}j-i&j-i-2&...&j-i-2k\\ (j-i)ji&(j-2-i)(i+1)(j-1)&...&(j-i-2k)(j-k)(i+k)\\ .&.&.&.\\ .&.&.&.\\ .&.&.&.\\ (j-i)ji...(j-(k+1-2))(i-(k+1-2))&&...&(j-i-2k)(j-k)(i+k)...(j-2k+1)(i+1)\end{pmatrix}$
where $i$ is the minimum such that $i+j=l$ and where we mean that every entry in the $r$-th row, $11\leq r\leq k+$ 1 (corresponding to an index $(i,j)$) is zero whenever $i\leq r-1$.
The determinant of the above matrix is equal to the product of $(j-i)...(j-i-2k)$ by the determinant of the following matrix
$\begin{pmatrix}1&1&...&1\\ ji&(i+1)(j-1)&...&(j-k)(i+k)\\ .&.&.&.\\ .&.&.&.\\ .&.&.&.\\ (j-i)ji...(j-(k+1-2))(i-(k+1-2))&&...&(j-k)(i+k)...(j-2k+1)(i+1)\end{pmatrix}$
Now consider the matrix above and subtract each column to the previous one. Then one gets a matrix
$\begin{pmatrix}1&0&...&0\\ ji&(j-1)-i&...&(j-1)-i-2(k-1)\\ .&.&.&.\\ .&.&.&.\\ .&.&.&.\\ (j-i)ji...(j-(k+1-2))(i-(k+1-2))&&...&(j-1-i-2(k-1))(j-1-(k-1))(i+k-1)...(j-2(k-1)+1)(i+1)\end{pmatrix}.$
Observe that
$\begin{pmatrix}(j-1)-i&...&(j-1)-i-2(k-1)\\ .&.&.\\ .&.&.\\ .&.&.\\ (j-1-i)(j-1)i...((j-1-(k-2)))(i-((k-2)))&...&(j-1-i-2(k-1))(j-1-(k-1))(i+k-1)...(j-2(k-1)+1)(i+1)\end{pmatrix}.$
is $B_{k,l-1}$ (where as we have already said, we mean that every entry corresponding to a index $(i,j)$ in the $r$-th row, $1\leq r\leq k$ is zero whenever $i\leq r-1$. The matrix has not zero determinant by the inductive hypothesis. Hence we have shown, by induction, that for any $33\leq l\leq 2g-$ 3 and for any $k\geq 2$, $(w_{l})^{2}$,…,$(w_{l})^{2k}$ impose exactly $c_{k,l}$ linearly independent conditions.
Now let $s_{k}$ be the number of the indexes $l$ such that $c_{k,l}>c_{k-1,l}$. Observe that the condition $c_{k,l}>c_{k-1,l}$ is equivalent to say that $l$-th equations of $Ker(\mu_{2k})$ are independent from the $l$-th equations of $Ker(\mu_{2}),...,Ker(\mu_{2(k-1)})$. Then we have
(41) $dim(Ker(\mu_{2k})=dimKer(\mu_{2(k-1)})-s_{k}.$
Notice that $c_{k,l}>c_{k-1,l}$ if and only if $k-1<n_{l}$. We claim that this happens if and only if $2k+1\leq l\leq 2g-(2k+1).$
In fact, if $2k+1\leq l\leq 2g-(2k+1)$, we have at least $k$ coordinates $(i,j)$ such that $i+j=l$. Indeed write $l$ as $l=2k+1+m$, $0\leq m\leq 2g-2(2k+1)$ (where we are using the assumption $2g-2(2k+1)\geq 0$).
If $l$ is odd (equivalently $m$ is even), these are given by
$(i,j)=(1+\frac{m}{2},2k+\frac{m}{2}),...,(k+\frac{m}{2},k+1+\frac{m}{2})$
These are easily seen to be admissible indexes $(i,j)$, that is they satisfy $1\leq i<j$, $i+j=l$ and $j\leq g-1$. If $l$ is even (equivalently $m$ is odd), these are given by
$(i,j)=(1+\lfloor\frac{m}{2}\rfloor,2k+\lfloor\frac{m}{2})\rfloor,...,(k+\lfloor\frac{m}{2}\rfloor,k+2+\lfloor\frac{m}{2}\rfloor).$
This shows that $n_{l}\geq k$.
On the other hand, if $l\leq 2k$ or $l\geq 2g-2k$, it is easy to see that the number of admissible indexes $(i,j)$ such that $i+j=l$ is strictly less than $k$. We then conclude that $s_{k}=2g-(2k+1)-2k=2g-(4k+1)$ and hence $dim(Ker(\mu_{2k})=dimKer(\mu_{2(k-1)})-s_{k}=dimKer(\mu_{2(k-1)})-(2g-(4k+1))$. So we have proven that $Rank(\mu_{2k})=2g-(4k+1)$.
By the equality: $dim(Ker(\mu_{2k}))=dimKer(\mu_{2(k-1)})-(2g-(4k+1))$, and using that $dim(I_{2})=\frac{(g-1)(g-2)}{2}$, we get
$dim(Ker(\mu_{2k}))=\frac{(g-1)(g-2)}{2}-\sum_{i=1}^{k}(2g-(4i+1))=\frac{(g-1)(g-2)}{2}+k(2k-2g+3),$
for every $k\leq\frac{g-1}{2}$. So if $g$ is odd and $k=\frac{g-1}{2}$, we have $dim(Ker(\mu_{2k}))=dim(Ker(\mu_{g-1}))=0$, hence for $k>\frac{g-1}{2}$, $\mu_{2k}\equiv 0$. If $g$ is even, and $k=\frac{g-2}{2}$, we have $dim(Ker(\mu_{2k}))=dim(Ker(\mu_{g-2}))=0$, hence for $k>\frac{g-2}{2}$, $\mu_{2k}\equiv 0$.
4. Schiffer variations and Weierstrass points
In this section we recall the definition and the basic properties of (higher) Schiffer variations.
Let $C$ be a smooth curve of genus $g>1$, take a point $p\in C$ and fix a local coordinate $z$ centred in $p$. For $31\leq n\leq 3g-$ 3, we define the $n^{th}$ Schiffer variation at $p$ to be the element $\xi_{p}^{n}\in H^{1}(C,T_{C})\cong H^{0,1}_{\bar{\partial}}(T_{C})$ whose Dolbeault representative is given by ¯$\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}$, where $\rho_{p}$ is a bump function in $p$ which is equal to one in a small neighborhood $U$ containing $p$, $\xi_{p}^{n}=[\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}]$. Clearly $\xi_{p}^{n}$ depends on the choice of the local coordinate $z$. Take $31\leq n\leq 3g-$ 3. Consider the exact sequence
$0\rightarrow T_{C}\rightarrow T_{C}(np)\rightarrow T_{C}(np)_{|np}\rightarrow 0,$
and the induced exact sequence in cohomology:
$0\rightarrow H^{0}(T_{C}(np))\rightarrow H^{0}(T_{C}(np)_{|np})\stackrel{{\scriptstyle\delta^{n}_{p}}}{{\rightarrow}}H^{1}(T_{C}).$
By Riemann Roch, if $n<2g-2$, or $n\leq 3g-3$ and $p$ is a general point, we have:
$h^{0}(T_{C}(np))=0.$
Hence we have an inclusion
$\delta^{n}_{p}:H^{0}(T_{C}(np)_{|np})\cong{\mathbb{C}}^{n}\hookrightarrow H^{1}(T_{C})$
and the image of $\delta^{n}_{p}$ in $H^{1}(C,T_{C})$ is the $n$-dimensional subspace generated by $\xi_{p}^{1},...,\xi_{p}^{n}$ (see [9]) for more details).
Given an element $\zeta\in H^{1}(C,T_{C})$, consider the map given by cup product:
$\cup\zeta:H^{0}(C,K_{C})\rightarrow H^{1}(C,{\mathcal{O}}_{C}).$
The rank of $\zeta$ is by definition the rank of the map $\cup\zeta$.
Remark 4.1.
Notice that $H^{0}(K_{C}(-np))\subset Ker(\cup\xi_{p}^{n})$, for $n\leq g$. In fact, $\omega\in ker(\cup\xi_{p}^{n})$ if and only if for any $\alpha\in H^{0}(K_{C}),$ we have $\xi_{p}^{n}(\omega\alpha)=0$. So if $z$ is a local coordinate around $p$ and $\omega=f(z)dz$, $\alpha=g(z)dz$ are local expressions, by [9, Lemma 2.2] we have:
$\xi_{p}^{n}(\omega\alpha)=\frac{2\pi i}{(n-1)!}(fg)^{(n-1)}(p)=0,$
if $\omega\in H^{0}(K_{C}(-np))$, $\forall\alpha\in H^{0}(K_{C})$. Hence $\xi_{p}^{n}$ has rank $\leq n$.
Let $\mathcal{H}_{g}$ be the moduli space of hyperelliptic curves of genus $g$. If $C$ is a hyperelliptic curve, then the tangent space of $\mathcal{H}_{g}$ at $[C]$ is $H^{1}(T_{C})^{+}$, that is the invariant subspace with respect to the hyperelliptic involution $\sigma$.
Assume $p\in C$ is a Weierstrass point. From the Dolbeault representation of the Schiffer variations, one immediately sees that the elements $\xi_{p}^{2k+1}$ are $\sigma$-invariant, hence they belong to $H^{1}(T_{C})^{+}$. So we have a subspace
$\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{2g-3}\rangle\subset H^{1}(T_{C})^{+}$
of dimension $g-1$.
Denote as usual by $|L|$ the $g^{1}_{2}$. We conclude this section giving a basis of $H^{0}(K_{C}\otimes L^{\vee})$ that will be useful in the next sections.
Remark 4.2.
Consider the hyperelliptic curve $C$ with equation (29). Fix the Weierstrass point $p=(0,0)$. For the computations of the next section it is more convenient to use the basis of $H^{0}(M)$ given by $\omega_{i}:=\alpha_{g-i}=\frac{x^{g-i}dx}{y}$, $i=1,...,g-1$. Notice that we have
$ord_{p}\omega_{k}=2g-2k-2,\ for\ all\ k=1,...,g-1.$
Hence we get:
  • $H^{0}(M)=\langle\omega_{1},...,\omega_{g-1}\rangle.$
  • $H^{0}(M(-p))=H^{0}(M(-2p))=\langle\omega_{1},...,\omega_{g-2}\rangle.$
    $\vdots$
  • $H^{0}(M(-(2g-4)p))=H^{0}(M(-(2g-5)p))=\langle\omega_{1}\rangle.$
So if we take a quadric $Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}(s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i})\in I_{2},$ as in Lemma 3.6 or Lemma 3.7, we have:
$Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}(s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i})=-\sum\limits_{1\leq k<h\leq g-1}a_{g-h,g-k}(s\omega_{k}\odot t\omega_{h}-s\omega_{h}\odot t\omega_{k})=$
$=\sum\limits_{1\leq k<h\leq g-1}b_{k,h}(s\omega_{k}\odot t\omega_{h}-s\omega_{h}\odot t\omega_{k}),$
with $b_{k,h}=-a_{g-h,g-k}$.
5. Second fundamental form of the Hyperelliptic Torelli locus
Let $C$ be a hyperelliptic curve of genus $g\geq 3$, $L$ be the line bundle giving the $g^{1}_{2}$. Set $M=K_{C}\otimes L^{\vee}$ and $\pi:C\rightarrow{\mathbb{P}}^{1}$ the map induced by $|L|$.
Denote by $\sigma$ the hyperelliptic involution and consider the decomposition $H^{0}(C,K_{C}^{\otimes 2})\cong H^{0}(C,K_{C}^{\otimes 2})^{+}\oplus H^{0}(C,K_{C}^{\otimes 2})^{-}$ of $H^{0}(C,K_{C}^{\otimes 2})$ in invariant and anti-invariant subspaces under the action of $\sigma$.
Denote by ${\mathcal{H}}_{g}$ the hyperelliptic locus in ${\mathcal{M}}_{g}$ and by
$j_{h}:{\mathcal{H}}_{g}\rightarrow{\mathcal{A}}_{g}$
the restriction of the Torelli map to ${\mathcal{H}}_{g}$. We endow ${\mathcal{A}}_{g}$ with the Siegel metric, that is the orbifold metric induced by the symmetric metric on the Siegel space $Sp(2g,{\mathbb{R}})/U(g)$ of which ${\mathcal{A}}_{g}$ is the quotient under the action of $Sp(2g,{\mathbb{Z}})$. The map $j_{h}$ is an orbifold immersion (see [15]) and we have the following tangent bundle exact sequences
[…]
Denote by
(43) $\rho_{HE}:N^{*}_{\mathcal{H}_{g}|\mathcal{A}_{g}}\rightarrow Sym^{2}\Omega^{1}_{\mathcal{H}_{g}},$
the dual of the second fundamental form of $j_{h}$.
At a point $[C]\in{\mathcal{H}}_{g}$, the dual of (42) is
[…]
where $m$ is the multiplication map and $I_{2}$ can be identified with the vector space of quadrics containing the rational normal curve.
So at a point $[C]$ the map $\rho_{HE}$ is a linear map
$I_{2}\rightarrow Sym^{2}(H^{0}(C,K_{C}^{\otimes 2})^{+}).$
In [5, Prop. 5.1] it is proven that $\forall Q\in I_{2}$, $\forall v,w\in H^{1}(T_{C})^{+}$ we have
(44) $\rho_{HE}(Q)(v\odot w)=\rho(Q)(v\odot w),$
where $\rho$ is the Hodge Gaussian map introduced in [8, Proposition-Definition 1.3].
In [11] the second fundamental form $\rho_{HE}$ of the hyperelliptic locus has been studied using the Hodge Gaussian map $\rho$ and the second gaussian map of the canonical bundle. In [11, Theorem 6.2] it has been proven that if $Y$ is a germ of totally geodesic subvariety of ${\mathcal{A}}_{g}$ generically contained in the hyperelliptic Torelli locus, then its dimension is at most $g+1$.
In [9] a relation between higher even Gaussian maps of the canonical bundle on a smooth projective curve of genus $g\geq 4$ and the second fundamental form of the Torelli map is given. This is a generalisation of a result obtained by Colombo, Pirola and Tortora on the second Gaussian map and the second fundamental form in [8] (see also [6, Theorem 2.2]). In fact, in [6, Theorem 2.2]) it is proven that for any $p\in C$, and for any $Q\in I_{2}$, we have
(45) $\rho(Q)(\xi_{p}\odot\xi_{p})=-2\pi i\mu_{2}(Q)(p).$
More precisely, the computations of $\rho(Q)$ on Schiffer variations given in [9, Proposition 3.1 and Remark 3.2], for quadrics $Q\in Ker(\mu_{2k})$ will be the main tool in the next section.
6. Isotropic subspaces
In this section, given a hyperelliptic curve $C$ of genus $g\geq 3$ and a Weierstrass point $p\in C$, we will determine for any $k\leq\lfloor\frac{g-3}{2}\rfloor$, the maximum $m\leq\lfloor\frac{g-1}{2}\rfloor$, such that the subspace $V_{m}=\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{2m+1}\rangle\subset V_{\lfloor\frac{g-1}{2}\rfloor}\subset H^{1}(T_{C})^{+}$ is isotropic with respect to all the quadrics $\rho(Q)$, $\forall Q\in Ker(\mu_{2k})$. We will show that $m=k$, namely that $V_{k}$ is isotropic for $\rho(Q)$, $\forall Q\in Ker(\mu_{2k})$, while $V_{k+1}$ is not (see Theorems 6.5, 6.7).
In order to do this we recall that, by (44), it suffices to compute $\rho(Q)(\xi_{p}^{r}\odot\xi_{p}^{n})$, for odd $r,n$.
To do this we will use [9, Remark 3.2], that we now recall for the reader’s convenience. Assume $Q$ is quadric in $I_{2}$, $Q=\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta},$ where $\gamma_{1},...,\gamma_{u}$ are elements of $H^{0}(K_{C})$. Choose a local coordinate $z$ around $p$ and take a local expression of $\gamma_{\alpha}$ around $p$: $\gamma_{\alpha}=g_{\alpha}(z)dz$.
Remark 6.1.
([9, Remark 3.2]) Assume
$\sum_{\alpha,\beta}c_{\alpha\beta}g_{\alpha}^{(h)}(0)g_{\beta}^{(l)}(0)=0,\ \forall h,l\geq 0,\ h+l\leq m,$
then
$\rho_{HE}(Q)(\xi_{p}^{n}\odot\xi_{p}^{r})=\rho(Q)(\xi_{p}^{n}\odot\xi_{p}^{r})=0,\ if\ r+n\leq m,\ \forall\ \text{odd}\ r,n\geq 1,$
and if $r+n=m+1$, then
$\rho_{HE}(Q)(\xi_{p}^{n}\odot\xi_{p}^{r})=\rho(Q)(\xi^{n}_{p}\odot\xi^{r}_{p})=2\pi i\left(\sum_{k=0}^{n-1}\left(\sum^{u}_{\alpha,\beta=1}c_{\alpha\beta}g^{(m+1-k)}_{\alpha}(0)g_{\beta}^{(k)}(0)\right)\frac{(n-k)}{k!(m+1-k)!}\right).$
Recall that a basis of $I_{2}$ is given by the following quadrics (see Lemma 3.3):
(46) $Q_{ij}:=s\omega_{i}\odot t\omega_{j}-s\omega_{j}\odot t\omega_{i}$
for $11\leq i<j\leq g-$ 1, where $L$ is the $g^{1}_{2}$, $H^{0}(L)=\langle s,t\rangle$, and $H^{0}(M)=H^{0}(K_{C}\otimes L^{\vee})=\langle\omega_{1},...,\omega_{g-1}\rangle$, where the sections $\omega_{i}$ are as in Remark 4.2. Let $Q$ be a quadric in $I_{2}$, then
$Q=\sum\limits_{1\leq i<j\leq g-1}b_{ij}Q_{ij}=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta},$
where $\gamma_{1},...,\gamma_{u}$ are elements of $H^{0}(K_{C})$. Choose a local coordinate $z$ around $p$ and take a local expression of $\gamma_{\alpha}$ around $p$: $\gamma_{\alpha}=g_{\alpha}(z)dz$.
Now let $k\geq 0$ and assume $Q\in Ker(\mu_{2k})$. Recall from Section 2 that
$Q\in Ker(\mu_{2k})\Leftrightarrow\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}\equiv 0\quad\forall h+l\leq 2k+1.$
Moreover by Lemma 3.5 we have
$\mu_{2k+2}(Q)=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}=\mu_{1,L}(s\wedge t)\mu_{2k+1,M}(\sum_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}\wedge\omega_{j}))=$
$=(s^{\prime}t-t^{\prime}s)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(r)}\omega_{j}^{(s)}-\omega_{i}^{(s)}\omega_{j}^{(r))})(dz)^{2k+4}\quad\forall h+l=2k+2,\quad\forall r+s=2k+1.$
It follows that
$\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=0\quad\forall h+l=2k+2,$
because $(s^{\prime}t-t^{\prime}s)(p)=0$, since $p$ is a Weierstrass point.
Now we want to determine the least $m\geq 2k+3$ such that
$\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)\neq 0,$
for some $h,l$ with $h+l=m.$
Lemma 6.2.
Let $Q=\sum\limits_{1\leq r<m\leq g-1}b_{r,m}Q_{r,m}\in I_{2}$, where the quadrics $Q_{r,m}$ are considered with respect to the basis $\{\omega_{i}\}_{i=1,...,g-1}$ given in Remark 4.2. If $Q\in Ker(\mu_{2k})$, we have
$b_{r,m}=0\quad\forall\quad r+m\geq 2g-(2k+2)\qquad(hence\ m\geq g-k).$
Proof.
Recall that we have $Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}(s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i})$, where $a_{ij}=-b_{g-j,g-i}$ (see Remark 4.2). We will prove by induction on $k$ that
(47) $a_{ij}=0\quad\forall\quad i+j\leq 2k+2.$
Notice that this finishes the proof. Indeed, putting $r=g-j,m=g-i$ we have
$i+j\leq 2k+2\Leftrightarrow r+m=2g-(i+j)\geq 2g-(2k+2).$
We also get $m\geq g-k$, because if $m\leq g-k-1$, $r<m$ implies $r+m<2g-2k-2$.
Recall that in Lemma 3.6 we have shown that $Q\in Ker(\mu_{2})$ if and only if for every $33\leq l\leq 2g-$ 3,
(48) $\sum\limits_{\begin{subarray}{c}1\leq i<j\\ i+j=l\end{subarray}}a_{ij}(j-i)=0.$
By Lemma 3.7, $Q\in Ker(\mu_{2k})$, $k\geq 2$, if and only if $Q\in Ker(\mu_{2})$ and and for every $m,l$ such that $2\leq m\leq k$, $32m-1\leq l\leq 2g-$ 3,
(49) $\sum\limits_{\begin{subarray}{c}1\leq i<j\\ i+j=l\\ i\geq m-1,\\ j\geq m-1\end{subarray}}a_{ij}(j-i)ij(i-1)(j-1)...(i-(m-2))(j-(m-2))=0.$
So let us prove (47). The case $k=1$ is trivial, since $42k+2=$ 4 and we only have $l=3,4$. From equations (48) we have $a_{12}=a_{13}=0$. Now assume that the result is true for $k-1$, where $k\geq 2$. Hence if $Q\in Ker(\mu_{2(k-1)})$, we have $a_{ij}=0\quad\forall l=i+j\leq 2k$. Now if $Q\in Ker(\mu_{2k})$ then $Q\in Ker(\mu_{2k-2})$, hence $a_{ij}=0\quad\forall l\leq 2k$. So we only have to check the cases $l=2k+1,2k+2$.
  • If $l=2k+1$, then we claim that $i\leq k$. Otherwise, if $i\geq k+1$, we have $j>i\geq k+1$, hence $i+j\geq 2k+3$, a contradiction. In the equations (49) with $l=2k+1$, we have $a_{1,2k},a_{2,2k-1},...,a_{k,k+1}$ as variables, since $i\leq k$. The equations given by (48) and (49) for $m=2,...,k$, $l=2k+1$, are then a system of $k$ equations in $k$ variables, which are linearly independent by the proof of Theorem 3.1. So we get $a_{i,j}=0$ for $i+j=2k+1$.
  • If $l=2k+2$, we get $i\leq k$ exactly as in the previous case. Hence again in (49) with $l=2k+2$, we have exactly $k$ linearly independent equations in the $k$ variables $a_{1,2k+1},a_{2,2k},...,a_{k,k+2}$. So $a_{ij}=0$ if $i+j=2k+2$.
Remark 6.3.
Let $\omega_{1},...,\omega_{g-1}$ be our basis of $H^{0}(M)$ and let $z$ be a local coordinate centered in the Weierstrass point $p$. Assume that we locally have $\omega_{i}=g_{i}(z)dz\otimes l^{-1}$, by abuse of notation we will write $\omega_{i}^{(h)}(p)$ instead of $g_{i}^{(h)}(0)$. Then by Remark 4.2 we have
$\omega_{i}^{(h)}(p)=0\quad\forall h\leq 2g-(2i+3).$
Lemma 6.4.
With the above notation, we have
$\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=0\quad\forall h+l\leq 4k+1.$
Proof.
We may assume $h>l$, since if $h=l$ we have trivially zero. From Remark 6.3, $\omega_{j}^{(l)}(p)=0\quad\forall j$ such that $2j+3\leq 2g-l$. Moreover, since $i<j$, this also implies $\omega_{i}^{(l)}(p)=0$, so the terms in the sum with $2j+3\leq 2g-l$ are identically zero. Then we can assume that the indexes $j$ that appear in the sum are those satisfying $12j+3\geq 2g-l+$ 1, which gives $j\geq g-\frac{l+2}{2}$. Denote by $\tilde{l}:=\frac{l+2}{2}$. Hence we have
$\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=\sum\limits_{1\leq i<j\leq g-1,j\geq g-\tilde{l}}b_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l)}-\omega_{i}^{(l)}\omega_{j}^{(h)})(p)=$
$=\sum\limits_{1\leq i<g-1}b_{i,g-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p)+\sum\limits_{1\leq i<g-2}b_{i,g-2}(\omega_{i}^{(h)}\omega_{g-2}^{(l)}-\omega_{i}^{(l)}\omega_{g-2}^{(h)})(p)\\ +\dots+\sum\limits_{1\leq i<g-\tilde{l}}b_{i,g-\tilde{l}}(\omega_{i}^{(h)}\omega_{g-\tilde{l}}^{(l)}-\omega_{i}^{(l)}\omega_{g-\tilde{l}}^{(h)})(p).$
Let us focus on the first term of the above sum. From Lemma 6.2, $Q\in Ker(\mu_{2k})$ implies $b_{i,g-1}=0\quad\forall g-(2k+1)\leq i\leq g-2$. Then we have
(50) $\sum\limits_{1\leq i<g-1}b_{i,g-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p)=\sum\limits_{i=1}^{g-2k-2}b_{i,g-1}(\omega_{i}^{(h)}\omega_{g-1}^{(l)}-\omega_{i}^{(l)}\omega_{g-1}^{(h)})(p).$
Moreover, again by Remark 6.3, if $2i+3\leq 2g-h<2g-l$, then $\omega_{i}^{(h)}(p)=\omega_{i}^{(l)}(p)=0$. But in the above sum we have $12i+3\leq 2(g-2k-2)+3=2g-4k-$ 1. Hence, if $h\leq 4k+1$ (so $l<4k+1$), then $2g-4k-1\leq 2g-h<2g-l$, so $2i+3\leq 2g-h<2g-l$ in all summands in (50), so the sum is identically zero.
Let us now consider the general term, where $j=g-s$. From Lemma 6.2, we have $b_{i,g-s}=0$ $\forall g-(2k+2-s)\leq i<g-s$. Hence the $i$’s that actually appear in the sum are those satisfying $i\leq g-(2k+2-s)-1=g-2k-3+s$. In the same way as we did for the first term, for $h\leq 4k+3-2s$, we have $2i+3\leq 2(g-2k-3+s)\leq 2g-h$, so $\omega_{i}^{(h)}(p)=\omega_{i}^{(l)}(p)=0$ and the term is zero.
It remains to look at the cases when $h\geq 4k+4-2s$. Since $h+l\leq 4k+1$ by assumption, we must have $14k+4-2s\leq h+l\leq 4k+$ 1, which gives $l\leq 2s-3$. For these values of $l$, we have $\omega_{g-s}^{(l)}(p)=0$ by Remark 6.3. Hence the term is zero again, because $i<j$, so $\omega_{i}^{(l)}(p)=0$. Thus we have shown that all terms are zero. ∎
We are now ready to state and prove our result.
Theorem 6.5.
Let $C$ be an hyperelliptic curve of genus $g$, $p\in C$ a Weierstrass point. Let $\gamma_{1},...,\gamma_{r}\in H^{0}(K_{C})$ where locally $\gamma_{\alpha}=g_{\alpha}dz$ and let
$Q=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta}=\sum\limits_{1\leq i<j\leq g-1}b_{ij}Q_{ij}\in Ker(\mu_{2k}).$
Then
$\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=0\quad\forall h+l\leq 4k+3.$
Proof.
We have to compute the following expression $\forall h+l\leq 4k+3$:
(51) $\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=\\ =\sum\limits_{1\leq i<j\leq g-1}b_{ij}[(s\omega_{i})^{(h)}(t\omega_{j})^{(l)}+(t\omega_{j})^{(h)}(s\omega_{i})^{(l)}-(s\omega_{j})^{(h)}(t\omega_{i})^{(l)}-(t\omega_{i})^{(h)}(s\omega_{j})^{(l)}](p).$
We have
$(s\omega_{i})^{(h)}(t\omega_{j})^{(l)}-(s\omega_{j})^{(h)}(t\omega_{i})^{(l)}=\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}s^{(h-n)}t^{(l-m)}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)}).$
Hence the term inside the square brackets of Equation 51 becomes
$\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)}).$
So equation (51) is
(52) $\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p).$
From Lemma 6.4 we have that
$\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)=0\quad\forall m+n\leq 4k+1.$
Since $n\leq h,m\leq l$ we have $m+n\leq h+l$, so this implies that the expression in (52) is zero for all $h,l$ such that $h+l\leq 4k+1$.
Now assume $h+l=4k+2$. If $n+m\leq 4k+1$ the expression is zero exactly as before, so the only case to consider is $n+m=4k+2$ and we must have $n=h,m=l$. But in this case we have trivially zero, because $(st-ts)=0$.
We are only left to the case $h+l=4k+3$. Now again, if $m+n\leq 4k+1$ we get zero and if $m+n=4k+3$ we get zero as in the previous case (this would imply again $n=h,m=l$). So we can only have $n+m=4k+2$. This implies either $n=h-1,m=l$ or $n=h,m=l-1$. In both cases, the first factor of (52) is (up to scalar) $(s^{\prime}t-t^{\prime}s)(p)=0$ since $p$ is a Weierstrass point. Hence we have proven that (52) is zero $\forall h+l\leq 4k+3$ and this concludes the proof.
For any $k\leq 2g-3$, consider the $(k+1)$-dimensional subspace $V_{k}:=\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{2k+1}\rangle\subset H^{1}(T_{C})^{+}$. We have the following
Theorem 6.6.
Let $0\leq k\leq\lfloor\frac{g-3}{2}\rfloor$ and let $Q\in Ker(\mu_{2k})$. Then we have
$\rho_{HE}(Q)(\xi_{p}^{l}\odot\xi_{p}^{m})=\rho(Q)(\xi_{p}^{l}\odot\xi_{p}^{m})=0\quad\forall l+m\leq 4k+3.$
This implies that the subspace $V_{k}=\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{2k+1}\rangle\subset H^{1}(T_{C})^{+}$ is isotropic for $\rho_{HE}(Q)$, $\forall Q\in Ker(\mu_{2k}).$
Proof.
The second assertion follows directly from the first one. The first one follows immediately by Remark 6.1 ([9, Remark 3.2]) and by Theorem 6.5. ∎
We have the following theorem.
Theorem 6.7.
Assume $0\leq k\leq\lfloor\frac{g-3}{2}\rfloor$. Then there exists $Q\in Ker(\mu_{2k})$ such that $\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+3})\neq 0$. Hence the subspace $V_{k+1}=\langle\xi_{p}^{1},...,\xi_{p}^{2k+1},\xi_{p}^{2k+3}\rangle\subset H^{1}(T_{C})^{+}$ is not isotropic for $\rho_{HE}(Ker(\mu_{2k}))$.
Proof.
Let $Q=\sum\limits_{\alpha,\beta=1}^{r}c_{\alpha,\beta}\gamma_{\alpha}\gamma_{\beta}$, where $\gamma_{1},...,\gamma_{r}\in H^{0}(K_{C})$. We will prove that there exists $Q$ such that $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})\neq 0$. By Remark 6.1, we know that
(53) $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=2\pi i\sum\limits_{u=0}^{2k}\left[\sum\limits_{\alpha,\beta=1}^{r}c_{\alpha,\beta}\gamma_{\alpha}^{(4k+4-u)}(0)\gamma_{\beta}^{(u)}(0)\right]\frac{(2k+1-u)}{u!(4k+4-u)!}$
Let us focus our attention on the term inside the square brackets, which is
(54) $\sum\limits_{\alpha,\beta=1}^{r}c_{\alpha,\beta}\gamma_{\alpha}^{(4k+4-u)}(0)\gamma_{\beta}^{(u)}(0).\quad$
We see that we must compute the derivatives of order $(h,l)$ of $Q$ where $h+l=4k+4$ and $0\leq l\leq 2k$. In order to do this, we will again use our basis of $H^{0}(K_{C}\otimes L^{\vee})$ given by the $\omega_{i}$’s, exactly as in (51) and (52). Hence we must compute
$\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)$
where $h+l=4k+4$. If $n+m\leq 4k+1$, we know that it is zero by Lemma 6.4. If $n+m=4k+4$, then in the equation above, we have $n=h$, $m=l$, so the first factor is (up to scalar) $(st-ts)=0$. If $n+m=4k+3$, we have either $n=h-1$ and $m=l$ or $n=h$ and $m=l-1$. In both cases, the first factor is (up to scalar) $(s^{\prime}t-t^{\prime}s)(p)=0$, because $p$ is a Weierstrass point. So we can only have $n+m=4k+2$, where $h=n$ and $m=l-2$ or $h=n-2$ and $m=l$, since if $n=h-1$ and $m=l-1$, the first factor is $(s^{\prime}t^{\prime}-t^{\prime}s^{\prime})=0$.
Summarising, in the above equation we only have to consider the following terms:
  1. (1)
    $n=h$ and $m=l-2$, which only works if $l\geq 2$ and gives
    $(st^{\prime\prime}-ts^{\prime\prime})(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(h)}\omega_{j}^{(l-2)}-\omega_{j}^{(h)}\omega_{i}^{(l-2)})(p).$
  2. (2)
    $n=h-2$ and $m=l$, which gives
    $(s^{\prime\prime}t-t^{\prime\prime}s)(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(h-2)}\omega_{j}^{(l)}-\omega_{j}^{(h-2)}\omega_{i}^{(l)})(p).$
Let us consider the first case. By Remark 6.3, we have $\omega_{j}^{(l-2)}(p)=0$ if $22j+3\leq 2g-l+$ 2, which gives $12j\leq 2g-l-$ 1. We have two subcases here: $l$ odd or $l$ even.
  • Assume $l$ is odd. We claim that we get zero. In fact, if $j\leq g-\frac{l+1}{2}$, then $\omega_{j}^{(l-2)}(p)=0$ by Remark 6.3. Hence we can assume $j\geq g-\frac{l+1}{2}+1=g-\frac{l-1}{2}$. Let $j:=g-u$, hence $u\in\{1,...,\frac{l-1}{2}\}$. From Lemma 6.2 we have $b_{i,g-u}=0$ for $i\geq g-2k-2+u$, so we may assume $i\leq g-2k-3+u$.
    We know that $\omega_{i}^{(l-2)}(p)=0$ if $i\leq g-\frac{l+1}{2}$, but $i\leq g-2k-3+u\leq g-\frac{l+1}{2}$ if and only if $u\leq 3+2k-\frac{l+1}{2}$. Observe that this is true since $u\leq\frac{l-1}{2}$ and $\frac{l-1}{2}\leq 3+2k-\frac{l+1}{2}$, since $l\leq 2k$. Hence $\omega_{i}^{(l-2)}(p)=0$.
    Now let us consider $\omega_{i}^{(h)}(p)$. We know that $\omega_{i}^{(h)}(p)=0$ if $32i\leq 2g-h-$ 3, but $i\leq g-2k-3+u$ implies $2i\leq 2g-4k-6+2u$. Now, $32g-4k-6+2u\leq 2g-h-$ 3 if and only if $2u\leq 4k+3-h$, which gives $l-1\leq 4k+3-h$, which is true if $l+h=4k+4$, that is our assumption. Hence $\omega_{i}^{(h)}(p)=0$, for all $i$ and hence every term vanishes.
  • Now assume $l$ is even. If $j\leq g-\frac{l+1}{2}=g-\frac{l}{2}-\frac{1}{2}<g-\frac{l}{2}$, we get $\omega_{j}^{(l-2)}(p)=0$ by Remark 6.3. So we may assume $j\geq g-\frac{l}{2}$. Setting $j=g-u$, we get $u\leq\frac{l}{2}$. As in the odd case, we may assume $i\leq g-2k-3+u$. Now $\omega_{i}^{(l-2)}(p)=0$ if $i\leq g-\frac{l}{2}-1$ and $g-2k-3+u\leq g-\frac{l}{2}-1$ if and only if $u\leq 2k+2-\frac{l}{2}$. Since $u\leq\frac{l}{2}$, we get that the last inequality is true if and only if $l\leq 2k+2$, which is the case because $l\leq 2k$. Hence $\omega_{i}^{(l-2)}(p)=0$.
    We have that $\omega_{i}^{(h)}(p)=0$ if $i\leq g-\frac{h}{2}-\frac{3}{2}<g-\frac{h}{2}-1$, so $\omega_{i}^{(h)}(p)=0$ if $i\leq g-\frac{h}{2}-2$. Hence we can assume $i\geq g-\frac{h}{2}-1$, but $i\leq g-2k-3+u$ and $g-2k-3+u-(g-\frac{h}{2}-1)=u-2-2k+\frac{h}{2}\leq\frac{l}{2}+\frac{h}{2}-2-2k=0$ because $h+l=4k+4$. Hence if $u=l/2$, there is a unique nonzero term, which is precisely $b_{g-\frac{h}{2}-1,g-\frac{l}{2}}\omega_{g-\frac{h}{2}-1}^{(h)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)$.
Let us now consider the second case. Again, we should distinguish the cases $l$ odd and $l$ even. However, arguing exactly in the same way as in the previous case, we again get that if $l$ is odd every term is zero.
So let us assume $l$ even, then also $h$ is even. By Remark 6.3, $\omega_{j}^{(l-2)(p)}=0$ if $j\leq g-\frac{l}{2}-\frac{3}{2}<g-\frac{l}{2}-1$, so we may assume $j\geq g-\frac{l}{2}-1$. As usual we put $j=g-u$ and we have $u\leq\frac{l}{2}+1$. We have $\omega_{i}^{(l)}(p)=0$ if $32i\leq 2g-l-$ 3, but, as in the previous case, we can assume $i\leq g-2k-3+u$. Since $l\leq 2k$, and $u\leq\frac{l}{2}+1$, we have $32g-6-4k+2u\leq 2g-l-$ 3. So $32i\leq 2g-6-4k+2u\leq 2g-l-$ 3, and hence $\omega_{i}^{(l)}(p)=0$.
Again, by Remark 6.3, $\omega_{i}^{(h-2)}(p)=0$ if $12i\leq 2g-h-$ 1. As usual we may assume $2i\leq 2g-6-4k+2u$. We have $u\leq\frac{l}{2}+1$. We claim that the only nonzero term occurs for $u=\frac{l}{2}+1$ and $i=g-\frac{h}{2}$. In fact, if $u\leq\frac{l}{2}$, we have $12g-6-4k+2u\leq 2g-h-$ 1, so $\omega_{i}^{(h-2)}(p)=0$. So $u=\frac{l}{2}+1$, $i=g-\frac{h}{2}$ and the only term is $b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$
We have computed all the non-zero terms of (54). These are
  • $(st^{\prime\prime}-ts^{\prime\prime})(p)[\binom{l}{2}b_{g-\frac{h}{2}-1,g-\frac{l}{2}}\omega_{g-\frac{h}{2}-1}^{(h)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)-\binom{h}{2}b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)]$ if $l\geq 2$ and $h,l$ even;
  • $-(st^{\prime\prime}-ts^{\prime\prime})(p)\binom{h}{2}b_{g-\frac{h}{2},g-\frac{l}{2}-1}\omega_{g-\frac{h}{2}}^{(h-2)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)$ if $l=0$ and $h=4k+4$.
We are now ready to complete the proof. Since $h+l=4k+4$, this implies $\frac{h}{2}=2k+2-\frac{l}{2}$, so that $g-\frac{h}{2}=g-2-2k+\frac{l}{2}$. We have
$\rho(Q)(\xi_{p}^{2k+3},\xi_{p}^{2k+1})=2\pi i\sum_{\begin{subarray}{c}2\leq l\leq 2k\\ l\text{ even}\end{subarray}}\frac{(2k+1-l)}{l!(4k+4-l)!}(st^{\prime\prime}-ts^{\prime\prime})(p)\cdot\\ \left[\binom{l}{2}b_{g-3-2k+\frac{l}{2},g-\frac{l}{2}}\omega_{g-3-2k+\frac{l}{2}}^{(4k+4-l)}(p)\omega_{g-\frac{l}{2}}^{(l-2)}(p)-\binom{h}{2}b_{g-2-2k+\frac{l}{2},g-\frac{l}{2}-1}\omega_{g-2-2k+\frac{l}{2}}^{(4k+2-l)}(p)\omega_{g-\frac{l}{2}-1}^{(l)}(p)\right]-\\ -2\pi i(st^{\prime\prime}-ts^{\prime\prime})(p)\frac{2k+1}{(4k+4)!}\binom{4k+4}{2}b_{g-2k-2,g-1}\omega_{g-2k-2}^{(4k+2)}(p)\omega_{g-1}(p)$
Notice that every term in the equation is a linear combination of $b_{r,u}$ where $r+u=2g-2k-3$.
Let $\lambda_{k,u}$ be the coefficient of $b_{g-2k-3+u,g-u}$ where $u=1,...,k+1$. We will now see that $\lambda_{k,u}\neq 0$ for every $k,u$. For $1\leq u\leq k$, we have
$\lambda_{k,u}=\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)\left[\frac{2k+1-2u}{(4k+4-2u)!}\binom{2u}{2}-\binom{4k+6-2u}{2}\frac{2k+3-2u}{(4k+6-2u)!}\right]=\\ =\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)\frac{-8u^{3}+8u^{2}(k+1)-4ku-2k-3}{2(4k+4-2u)!}$
where the last equality follows after some easy computations. Notice that
$-8u^{3}+8u^{2}(k+1)-4ku-2k-3\neq 0$
because it is odd for every $k,u$. Moreover, $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)\neq 0$ because $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)$ has order $4k+4-2u$ at $p$ and $\omega_{g-u}^{(2u-2)}$ has order $22u-$ 2 at $p$ (see Lemma 4.2).
Finally, if $u=k+1$ we have
$\lambda_{k,k+1}=-\frac{1}{2(2k+2)!}\omega_{g-2-k}^{(2k+2)}(p)\omega_{g-k-1}^{(2k)}.$
Hence $\lambda_{k,u}\neq 0$, for all $u=1,...,k+1$.
So assume by contradiction that
(55) $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=\sum\limits_{u=1}^{k+1}\lambda_{k,u}b_{g-2k-3+u,g-u}=-\sum\limits_{u=1}^{k+1}\lambda_{k,u}a_{u,2k+3-u}=0$
for every $Q\in Ker(\mu_{2k})$, where $a_{ij}=-b_{g-j,g-i}$ by Remark 4.2. Recall that $Q\in Ker(\mu_{2k})$ if and only if the coefficients $a_{ij}$ satisfy the equations given in Lemma 3.7. Moreover observe that from Theorem 3.1 it follows that we have $k$ linearly independent equations in the variables $a_{u,2k+3-u}$ with $u=1,...,k+1$. We have shown that each coefficient $\lambda_{k,u}$ in equation (55) is different from zero and up to scalar it is the product $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)$. So we can multiply the sections $\omega_{g-u}$ by a non zero scalar multiple, hence equation (55) cannot be linearly dependent from the $k$ equations giving $Ker(\mu_{2k})$. Hence there exists a quadric $Q\in ker(\mu_{2k})$ such that $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})\neq 0$.
Remark 6.8.
Notice that we have shown that the equation (55) gives a hyperplane in $Ker(\mu_{2k})$ that we denote by
$A_{k,0}:=\{Q\in Ker(\mu_{2k}):\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})=0\}.$
Moreover we have just proven that a quadric $Q\in A_{k,0}$ iff satisfies the system of $k+1$ linearly independent equations in $k+1$ variables, given by the equations in Lemma 3.7 defining $Ker(\mu_{2k})$ and equation (55). In particular the coefficients of $Q$ satisfy the equations $b_{g-2k+3+u,g-u}=-a_{u,2k+3-u}=0\quad\forall u=1,...,k+1$.
Example 6.9.
If $k=0$ we have a very concrete description of $A_{0,0}$. In fact, we have $\rho(Q_{ij})(\xi_{p},\xi_{p}^{3})=0$, $\forall(i,j)\neq(g-2,g-1)$ and $\rho(Q_{g-2,g-1})(\xi_{p},\xi_{p}^{3})\neq 0$. Hence we get that $A_{0,0}=\langle Q_{i,j}\mid 1\leq i<j\leq g-1,(i,j)\neq(g-2,g-1)\rangle$.
Theorem 6.10.
Let $0\leq k\leq\lfloor\frac{g-3}{2}\rfloor$ and let us now assume $Q\in A_{k,0}$. Then there exist coefficients $\alpha_{k,u}$, all non-zero such that
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=\sum\limits_{u=1}^{k+1}\alpha_{k,u}b_{g-4-2k+u,g-u}\omega_{g-4-2k+u}^{(4k+6-2u)}(p)\omega_{g-u}^{(2u-2)}(p),$
Moreover,
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=0\Leftrightarrow b_{g-4-2k+u,g-u}=0\quad\forall u=1,...,k+1.$
Proof.
The proof is very similar to the one given in Theorem 6.7, where by Remark 6.1 we must now compute
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=2\pi i[\sum\limits_{u=0}^{2k+2}\sum\limits_{\alpha,\beta=1}^{r}c_{\alpha,\beta}\gamma_{\alpha}^{(4k+6-u)}(0)\gamma_{\beta}^{(s)}(0)]\frac{(2k+3-u)}{u!(4k+6-u)!}.$
By looking at the term inside the square brackets, we see that we must compute the derivatives of order $(h,l)$ of $Q$ where $h+l=4k+6$ and $20\leq l\leq 2k+$ 2. Hence we have to compute
(56) $\sum\limits_{n=0}^{h}\sum\limits_{m=0}^{l}\binom{h}{n}\binom{l}{m}(s^{(h-n)}t^{(l-m)}-t^{(h-n)}s^{(l-m)})(p)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{j}^{(n)}\omega_{i}^{(m)})(p)$
where $h+l=4k+6$.
The first step is analogous to Lemma 6.4: if $Q\in A_{k,0}$ we have
$\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(n)}\omega_{j}^{(m)}-\omega_{i}^{(m)}\omega_{j}^{(n)})(p)=0\quad\forall m+n\leq 4k+3.$
Then one immediately sees that in (56) we only have to consider the case $n+m=4k+4$. Now, after some computations which are analogous to the ones in Theorem 6.7, we get
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=\sum\limits_{u=1}^{k+1}\alpha_{k,u}b_{g-4-2k+u,g-u}\omega_{g-4-2k+u}^{(4k+6-2u)}(p)\omega_{g-u}^{(2u-2)}(p),$
where $\alpha_{k,u}\neq 0$ for every $u$. Then we conclude as in Theorem 6.7 that
$\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=0\Leftrightarrow b_{g-4-2k+u,g-u}=0\quad\forall u=1,...,k+1.$
Remark 6.11.
Let $Q\in A_{k,0}$. By Theorem 6.10, the linear condition $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=0$ gives a hyperplane in $A_{k,0}$. Let us denote by
$A_{k,0,0}:=\{Q\in A_{k,0}:\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+3})=0\}.$
Following [7] we give the following definition
Definition 6.12.
A nonzero direction $\zeta\in H^{1}(T_{C})^{+}$ is said to be asymptotic if $\rho_{HE}(Q)(\zeta\odot\zeta)=\rho(Q)(\zeta\odot\zeta)=0$, $\forall Q\in I_{2}$.
Notice that saying that a nonzero element $\zeta\in H^{1}(T_{C})^{+}$ is an asymptotic direction means that the point $[\zeta]\in{\mathbb{P}}H^{1}(T_{C})^{+}$ is in the base locus of the linear space of quadrics $\rho_{HE}(I_{2})\subset Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$.
Remark 6.13.
Recall that by Theorem 6.6 putting $k=0$, a Schiffer variation at a Weiestrass point $\xi_{p}$ is asymptotic. Moreover, moving a branch point in ${\mathbb{P}}^{1}$, one can see that there exist algebraic curves in the hyperelliptic locus having these Schiffer variations as tangent directions (see [12]).
Theorem 6.14.
Let $C$ be hyperelliptic of genus $g\geq 3$ and $p\in C$ be a Weierstrass point. Then the asymptotic directions in the space $V_{\lfloor\frac{g-2}{2}\rfloor}$, which is equal to $\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{g-1}\rangle$ if $g$ is even, and equal to $\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{g-2}\rangle$ if $g$ is odd, are exactly the Schiffer variations $\xi_{p}=\xi_{p}^{1}$.
Proof.
By Remark 6.13 we know that the Schiffer variations at the Weierstrass points are asymptotic.
To show that these are the only ones, we will prove the result when $g$ is even, the case $g$ odd being very similar.
Let $v\in\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{g-1}\rangle$, then $v=\sum\limits_{i=0}^{\frac{g-2}{2}}\lambda_{2i+1}\xi_{p}^{2i+1}$ and let $l$ be the maximum odd integer such $\lambda_{l}\neq 0$. We may assume $l\geq 3$, since $\rho(Q)(\xi_{p}\odot\xi_{p})=-2\pi i\mu_{2}(Q)(p)=0$ for all $Q\in I_{2}$. Then we have $v\in\langle\xi_{p}^{1},...,\xi_{p}^{l}\rangle$. Since $l$ is odd, we have $l=2k+1$ for some integer $1\leq k\leq\frac{g-2}{2}$. From Theorem 6.7 we know that there exists $Q\in Ker(\mu_{2k-2})$ such that $\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k-1})\neq 0$ and $\rho(Q)(\xi_{p}^{i}\odot\xi_{p}^{j})=0$ for every $i,j\leq 2k-1$, $i,j$ odd integers. Then we have
$\rho(Q)(v\odot v)=2\lambda_{2k-1}\lambda_{2k+1}\rho(Q)(\xi_{p}^{2k-1}\odot\xi_{p}^{2k+1})+$
$+\lambda_{2k+1}^{2}\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1}).$
First, notice that for every $k\leq\frac{g-2}{2}$ we have $A_{k-1,0}\subsetneq Ker_{\mu_{2k-2}}$ and $A_{k-1,0,0}\subsetneq A_{k-1,0}$. In fact, dim $Ker(\mu_{2k-2})\geq 3$, since dim $Ker(\mu_{g-4})=3$ by Theorem 3.1 and $42k-2\leq g-$ 4.
If $\lambda_{2k-1}=0$, we simply take $Q\in A_{k-1,0}\setminus A_{k-1,0,0}$ and we get $\rho(Q)(v\odot v)\neq 0$ as desired.
If $\lambda_{2k-1}\neq 0$, we choose $Q\in Ker(\mu_{2k-2})\setminus A_{k-1,0}$ and $Q^{\prime}\in A_{k-1,0}\setminus A_{k-1,0,0}$. Hence, for a linear combination $Q+aQ^{\prime}$, we get
$\rho(Q+aQ^{\prime})(v\odot v)=2\lambda_{2k-1}\lambda_{2k+1}\rho(Q)(\xi_{p}^{2k-1}\odot\xi_{p}^{2k+1})+$
$+\lambda_{2k+1}^{2}[\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})+a\rho(Q^{\prime})(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})].$
By the choice of $Q^{\prime}$, we have $\rho(Q^{\prime})(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})\neq 0$, so if we take
$a=-\frac{\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})}{\rho(Q^{\prime})(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})}$
we get $\rho(Q+aQ^{\prime})(v\odot v)=2\lambda_{2k-1}\lambda_{2k+1}\rho(Q)(\xi_{p}^{2k-1}\odot\xi_{p}^{2k+1})\neq 0$ since $Q\notin A_{k-1,0}$.
Remark 6.15.
In [5, Corollary 3.4] it is shown that $\rho_{HE}$ is injective, so
$\dim\rho_{HE}(Ker\mu_{2k})=\dim Ker\mu_{2k}=\frac{(g-1)(g-2)}{2}-k(2g-2k-3),$
for all $k$. For any hyperelliptic curve $C$ of genus $g\geq 3$ and for any Weiestrass point $p\in C$, we have proven in Theorems 6.6, 6.7 that we have the following chain of subspaces of quadrics in ${\mathbb{P}}H^{1}(T_{C})^{+}$ and of corresponding maximal isotropic subspaces in $V_{\lfloor\frac{g-1}{2}\rfloor}$.
If $g$ is odd:
$\rho_{HE}(Ker(\mu_{g-3}))\subsetneq\rho_{HE}(Ker(\mu_{g-5}))...\subsetneq\rho_{HE}(Ker(\mu_{2}))\subsetneq\rho_{HE}(I_{2});$
$V_{\frac{g-3}{2}}\supsetneq V_{\frac{g-5}{2}}\supsetneq...\supsetneq V_{1}\supsetneq V_{0}=\langle\xi_{p}\rangle,$
while if $g$ is even we have
$\rho_{HE}(Ker(\mu_{g-4}))\subsetneq\rho_{HE}(Ker(\mu_{g-6}))...\subsetneq\rho_{HE}(Ker(\mu_{2}))\subsetneq\rho_{HE}(I_{2}),$
$V_{\frac{g-4}{2}}\supsetneq V_{\frac{g-6}{2}}\supsetneq...\supsetneq V_{1}\supsetneq V_{0}=\langle\xi_{p}\rangle.$
Theorem 6.14 also allows us to give a bound for the dimension of a germ of a totally geodesic submanifold of ${\mathcal{A}}_{g}$, generically contained in the Hyperelliptic Torelli locus. Notice that, nevertheless this bound is weaker than the one proven in [11, Theorem 6.2].
Corollary 6.16.
Let $Y$ be a germ of a totally geodesic submanifold of $\mathcal{A}_{g}$ generically contained in $j(\mathcal{H}_{g})$ passing through $j(C)$, where $C$ is a hyperelliptic curve of genus $g$. Then we have
$\dim(Y)\leq\lfloor\frac{3g+1}{2}\rfloor.$
Proof.
We will prove it for $g$ even, the case $g$ odd is analogous.
Set $W:=T_{j(C)}Y\subset H_{1}(T_{C})^{+}$ and $V:=V_{\frac{g}{2}-1}=\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{g-1}\rangle$. We have dim$(V)=\frac{g}{2}$. We claim that $\dim(V\cap W)\leq 1$. Indeed, take $v\in V\cap W$, then $v$ is asymptotic, since $Y$ is totally geodesic. So, by Theorem 6.14 we know that $v$ is a multiple of $\xi^{1}_{p}$. Hence we have
[…]
so dim$(Y)$= dim $(W)\leq 2g-\frac{g}{2}=\frac{3}{2}g$. ∎